OpenStudy (hunt123):

suppose f(x)=x^2. find the graph of f(x+2)

jhonyy9 (jhonyy9):

like a first step find the f(x+2) = ?

OpenStudy (hunt123):

??

jhonyy9 (jhonyy9):

substitute the x+2 in place of x than you know that f(x)=x^2 so how will be f(x+2)=?

OpenStudy (hunt123):

im not to sure thats why im here

jhonyy9 (jhonyy9):

than f(x)=x^2 when substitute the (x+2) in place of x what will get ?

OpenStudy (joannablackwelder):

This form means to plug in x+2 for every time we see x in the function.

OpenStudy (joannablackwelder):

So, I first write the function without x's but leaving open parentheses in their place. ( )^2

OpenStudy (joannablackwelder):

Then plug in the x+2 into the parentheses. Can you try that step?

OpenStudy (mathstudent55):

This problem is simply a transformation. If you know what f(x) = x^2 looks like, then since you are asked for f(x + 2), you see that you are replacing x with x + 2. In a function, when you replace x with x - h, you end up with a translation of h units horizontally. In this case, f(x + 2) = f (x - (-2)) h = -2, and it is a translation of 2 units to the left.